# LoLA: Low-Rank Linear Attention With Sparse Caching

### v1

@@ 4 Experiments & Results
In our experiments, we leverage the same attention distillation procedure in LoLCATs to obtain the base model, then apply our inference strategy, LoLA, at test time. To train the baseline model, we replace each attention module in Llama-3.1 8B (or Llama-3.2 1B) with a hybrid sliding window + linear attention module. We use a sliding window size $\eta=64$ for training and use a trainable feature map for $\phi$ as described in (6). Here, the output dimension of $\phi$ is $D=2d$, where $d$ is the respective head dimension of the pretrained transformer (e.g., $d=128$ for Llama-3.1 8B). First, we freeze all non-attention layers in the linearized Transformer and only train $\phi$ with distillation for two epochs on Alpaca. Then, we perform LoRA [26] finetuning on the whole model for another two epochs. This procedure only uses 40M training tokens with 1024-long sequences. For the base LoLA used in Table 1 and 3, we keep LoLCATs’ sliding window size of $\eta=64$ and add a sparse cache of size $\lambda=64$. We also provide extended cache sizes in Table 1.
@@ 4.1 Associative Recall
To see how LoLA improves associative recall, we conduct a study on Single-Needle-in-a-Haystack (S-NIAH) tasks from RULER [25]. In these tasks, the “haystack” is synthetically constructed with various context lengths. The first task, S-NIAH-1, uses random sentences (e.g., “The grass is green.“) as the haystack, while S-NIAH-2 & 3 use essays. The needle—represented as a (word, number) pair—is placed in the haystack. At the end of the prompt, the model is tasked with returning the associated number with the special word. The first two tasks (S-NIAH-1 & 2) use a 7-digit number in the needle, and S-NIAH-3 uses a 32-digit UUID, requiring more tokens to represent the needle.
In our experiment, we ablated extended forms of LoLCATs-8B (i.e extended sliding window, no sparse caching) and LoLA-8B. We provided various total cache sizes, choosing $C=\lambda$ for simplicity. All extended caches are marked with “+”s. Each extension doubles the amount of tokens computed with softmax attention. For a simpler reference, we report the compression rate as the size of Llama-3.1-8B’s cache divided by the total cache size of each method. Since Llama’s cache depends on context length, we report multiple rates.
@@
Model Compression S-NIAH-1 S-NIAH-2 S-NIAH-3 Rate @ 2-4K .5K 1K 2K 4K .5K 1K 2K .5K 1K 2K Transformer Llama-3.1-8B 1$\times$ 100 100 100 100 100 100 100 100 100 100 Base Subquadratic Model LoLCATs-8B 11$\times$-22$\times$ 9.0 3.2 1.4 0.6 100 7.6 2.0 97.4 1.6 0.6 Extended at Inference LoLCATs-8B+ 6.4$\times$-13$\times$ 29.4 9.6 3.4 1.4 100 17.4 7.2 98.2 14.6 3.2 LoLA-8B 6.4$\times$-13$\times$ 99.0 95.4 79.4 69.4 100 39.4 3.0 99.8 7.4 1.6 LoLCATs-8B++ 4.0$\times$-8.0$\times$ 87.2 26.8 10.2 3.2 100 37.0 12.2 100 32.4 6.0 LoLA-8B++ 4.0$\times$-8.0$\times$ 100 99.6 96.4 89.6 100 98.8 15.0 99.8 33.4 9.2 LoLCATs-8B$+$ 2.3$\times$-4.6$\times$ 100 65.6 24.6 8.8 100 71.8 21.6 100 66.0 10.6 LoLA-8B$+$ 2.3$\times$-4.6$\times$ 100 100 99.6 97.4 100 100 85.4 99.8 99.8 27.2 LoLA-8B++++ 1.2$\times$-2.4$\times$ 100 100 100 99.0 100 100 100 100 100 100
In Table 1, we observe LoLCATs struggles to recall information outside the sliding window. Extending the sliding window size marginally improves the performance. For large haystacks, the accuracy is roughly the proportion of context that the sliding window covers. For smaller haystacks, the performance is slightly higher than that ratio since fewer pairs stored in the hidden state. This results in fewer collisions. Additionally, we observe that fewer collisions exist in essay-based haystacks (S-NIAH-2 & 3), likely as a result of being more similar to the distillation data.
Model .5K 1K 2K 4K 8K 16K LoLA-8B 4+ 100% / 1$\times$ 100% / 1$\times$ 100% / 1.2$\times$ 99.0% / 2.4$\times$ 92.2% / 4.9$\times$ 13.6% / 9.8$\times$
LoLA drastically improves recall with minimal additional caching. Table 1 demonstrates an improvement from 0.6% to 97.4% accuracy on S-NIAH-1 while still using a 4.6×4.6$\times$4.6 × smaller cache than Llama. This unlocks a new capability for hybrid linear attention architectures. Our extended results in Table 2 demonstrate recall on up to 8K context length, which is 8×8$\times$8 × longer than the sequences seen during training.
Extending the cache size can improve performance while still maintaining high compression rates over transformers. For even longer context lengths, LoLA’s cache can easily be scaled at inference to achieve the desired recall performance. LoLA is at least as memory-efficient as attention since we can always fall back to softmax attention for short sequences. Lastly, we observe that harder needle-in-a-haystack tasks (e.g., S-NIAH-3 with 32-digit needles) may require more sparse caching. To further extend these results, we believe training with longer sequences should yield stronger performance.
@@ 4.2 Commonsense Reasoning
We demonstrate the language modeling performance of LoLA on various zero-shot commonsense reasoning tasks using LM evaluation harness [17]. To compare against previously available approaches, we use PIQA (PI) [8], ARC-Easy (AE) & ARC-Challenge (AC) [12], HellaSwag (HS) [55], WinoGrande (WG) [40], MMLU (MM) [23], and Lambada OpenAI [33]).
In Table 3, we compare LoLA against other 7-9B subquadratic models (Mamba [21], Mamba2 [14], RWKV-6 [35], Hawk & Griffin [15], Falcon Mamba [58], RecurrentGemma [9], Mamba-in-the-Llama [49], Llamba [5], Hedgehog [57], and LoLCATs [56]). Models that use any form of unbounded global attention (e.g interleaving SSM blocks and Transformer blocks) still retain quadratic compute complexity and growing memory costs. These are outside the scope of this work. We also report the number of training tokens used to create each model in both tables. Though LoLA is an inference strategy that can be used for any sliding window + linear attention model, we report the cost of distilling the base subquadratic model [56]. In Appendix A, we show results for 1-2B subquadratic models. Additionally, we provide a direct comparison of distilled models by measuring the average accuracy relative to their teacher models.
Table 3: Performance comparison of 7-9B parameter fixed-memory models across various common sense reasoning tasks: PIQA (PI), Arc-Easy (AE), ARC-Challenge (AC), Winogrande (WG), MMLU (MM), and Lambada-openai (LB). Bolded scores are the best and underlined scores are the second best. We report accuracy for all applicable, except AC and HS use normalized logits. MMLU (MM) uses 5-shot. Reported scores were compiled from [5, 56, 48, 49]. * indicates our reproduced score is used and is higher than reported score.
Model Tokens (B) PI AE AC HS WG MM LB
@@
On short context tasks such as Winogrande, we observe additional caching is not needed as only local information is required for good performance. On the other hand, we find that gaps still exist with Lambada and MMLU. Lambada requires longer context reasoning, leading to significant improvements with sparse caching. Furthermore, Bick et al. [5] suggest that dataset selection plays a large role for MMLU performance. Though sparse caching shows significant improvement on MMLU, a more powerful distillation procedure or dataset may be needed to reach Llama’s performance [19].
@@ 4.3 Understanding Memory Collisions
We visualize how memory collisions occur in practice. At each time step $t$, we measure the self-recall error (9) for every KV pair that is currently stored in the hidden state $\mathbf{H}_{t}$. We visualize the error for linear attention, sliding window + linear attention, and LoLA. We use a sliding window size of $\eta=64$ tokens and a sparse cache size of $\lambda=64$ when applicable.
Figure 3: Visualizing memory collisions by measuring self-recall error for stored KV pairs.
When only using linear attention, we observe large recall errors. At early time steps, the first few KV pairs receive small errors, but quickly become larger after hidden state updates. In Appendix E, we additionally plot the relative error to better show how this occurs. These stored associative memories become corrupted and tough to recall in the future. Furthermore, difficult-to-memorize pairs are evident, illustrated as bright columns in Figure 3. The additional use of sliding window attention only delays the inevitable memory collisions.
@@
In our final experiment, we measure alternative scoring functions to understand which KV pairs should be sparsely cached. We provide an informal explanation with each function. Extended details for the score calculation can be found in Appendix D.
Table 4: Ablation results for various scoring methods on S-NIAH-1 with 512 context length.
Alternatives for Score$(\bm{k},\bm{v})$ S-NIAH-1 @ .5K Informal Assumption for “Important” Pairs $\left\|\cfrac{\phi(\bm{k})^{\top}\mathbf{H}}{\phi(\bm{k})^{\top}\bm{s}}-\bm{v}\right\|$ 99.0% Pairs that do not align with the hidden state’s prediction $(\exp(\bm{q}^{\top}\bm{k})-\phi(\bm{q})^{\top}\phi(\bm{k}))^{2}$ 11.4% Keys with incorrect attention weights $|\exp(\bm{q}^{\top}\bm{k})-\phi(\bm{q})^{\top}\phi(\bm{k})|$ 20.0% Keys with incorrect attention weights $\cfrac{\phi(\bm{q})^{\top}\phi(\bm{k})}{\exp(\bm{q}^{\top}\bm{k})}$ 52.0% Keys that Linear Attention over estimates None, extend sliding window 29.4% Most recent pairs
In Table 4, we observe that storing keys with poor exponential dot product approximations underperforms the naive extension of sliding window attention. This hints that a better softmax approximation should not be the main objective for linear attention methods. Keys over-estimated by linear attention seem to be “more important” than local keys; however, all tested alternatives fall short of enabling associative recall. Overall, our experiments suggest that key-value pairs that are unaligned with the hidden state should be carefully handled.
@@ 5 Conclusion
LoLA integrates linear attention with sparse caching to effectively mitigate memory collisions. By selectively retaining KV pairs that do not align with the current hidden state, LoLA enables passkey retrieval when the base model fails. Our experimental results demonstrate that targeted sparse caching substantially improves performance over naively increasing sliding window size. LoLA demonstrates strong language modeling performance over other 1B or 8B subquadratic models.
Limitations and future work. The sparse cache carries a small overhead compute cost of $\mathcal{O}(\lambda d)$ for scoring. For high-complexity, long-context tasks, we found that larger sparse cache sizes are needed to reduce interference in the hidden state. We believe that these limitations can addressed in the future with a better base subquadratic model. More advanced architectures like Gated DeltaNet [51] or Titans [3] use stronger forms of associative memory, which may lead to smaller caches.
@@ Appendix A Extended Language Modeling Results
Following Section 4.2, we extend this comparison in Table 5 for various 1-2B parameter subquadratic models [20, 28, 44, 38, 41, 51, 21, 14, 2, 35, 6, 56, 5, 39]. Compared to the 8B models, we observe that sparse caching is more important in the 1B parameter regime since there exist more memory collisions. This is a direct result of a smaller hidden state dimension. The size of $\mathbf{H}_{t}$ scales with the head dimension of the base model. Llama-3.2 1B’s head dimension is half that of Llama-3.1 8B.
@@
Subquadratic: Pretrained from scratch on FineWeb-Edu [51, 34]
@@
Subquadratic: Pretrained from scratch on various sources [6, 5]
@@
Subquadratic: Distilled from Phi-1.5-1.3B [6, 56]
@@
Subquadratic: Distilled from Llama-3.2-1B [5, 56]
@@
Llamba-1.3B 8 74.0* 69.5* 37.2 61.2 60.6 48.4 58.5 0.975×0.975$\times$0.975 ×
LoLCATs-1.3B 0.04 74.6 63.0 35.1 63.7 61.5 53.4 58.6 0.977×0.977$\times$0.977 ×
LoLA-1.3B (ours) 0.04 76.2 66.2 36.9 64.1 60.9 61.9 61.0 1.017×1.017$\times$1.017 ×
@@
Mamba2-Llama3-8B 20 76.8 74.1 48.0 70.8 58.6 43.2 61.9 0.837×0.837$\times$0.837 ×
@@
Llamba-8B 12 80.9 82.5 54.6 77.6 73.3 69.4 73.1 0.987×0.987$\times$0.987 ×
LoLCATs-8B 0.04 81.0 82.4 54.4 79.1 73.6 67.6 73.0 0.987×0.987$\times$0.987 ×
LoLA-8B 0.04 81.6 82.5 55.4 79.8 73.6 74.9 74.6 1.009×1.009$\times$1.009 ×
@@
Appendix B Related Work

### v2

@@ 4 Experiments & Results
In our experiments, we leverage the same attention distillation procedure in LoLCATs to obtain the base model, then apply our inference strategy, LoLA, at test time. To train the base model, we replace each attention module in Llama-3.1 8B (or Llama-3.2 1B) with a hybrid sliding window + linear attention module. We use a sliding window size $\eta=64$ for training and use a trainable feature map for $\phi$ as described in equation 6. The output dimension of $\phi$ is $D=2d_{k}$. First, we freeze all non-attention layers in the linearized Transformer and only train $\phi$ with distillation for two epochs on Alpaca (Taori et al., 2023). Then, we perform LoRA (Hu et al., 2022) finetuning on the whole model for two epochs. This procedure only uses 40M training tokens with 1024-long sequences.
@@ 4.1 Associative Recall
To see how LoLA improves associative recall, we conduct a study on Single-Needle-in-a-Haystack (S-NIAH) tasks from RULER (Hsieh et al., 2024). In Table 1, we compare LoLA to the base model, LoLCATs-8B, and variants with an extended sliding window for a fair comparison. We observe LoLCATs struggles to recall information outside the sliding window. Extending the sliding window size marginally improves performance. We explain the differences of each NIAH task in Appendix C and discuss these results more in depth.
Table 1: Measuring long context recall with Needle-in-a-Haystack (NIAH) tasks from the RULER benchmark. We report recall accuracy for each method across different context lengths (512, 1024, etc.) for each task.
@@
Model Cache Params S-NIAH-1 S-NIAH-2 S-NIAH-3 $(\eta,\lambda)$ .5K 1K 2K 4K .5K 1K 2K .5K 1K 2K Transformer Llama-3.1-8B ($\infty$, 0) 100 100 100 100 100 100 100 100 100 100 Base Subquadratic Model LoLCATs-8B (64, 0) 9.0 3.2 1.4 0.6 100 7.6 2.0 97.4 1.6 0.6 Extended at Inference LoLCATs-8B+ (128, 0) 29.4 9.6 3.4 1.4 100 17.4 7.2 98.2 14.6 3.2 LoLA-8B (64, 64) 99.0 95.4 79.4 69.4 100 39.4 3.0 99.8 7.4 1.6 LoLCATs-8B+ (512, 0) 100 65.6 24.6 8.8 100 71.8 21.6 100 66.0 10.6 LoLA-8B (256, 256) 100 100 99.6 97.4 100 100 85.4 99.8 99.8 27.2 LoLA-8B (512, 512) 100 100 100 99.9 100 100 100 100 100 100
Next in Table 2, we measure the performance of LoLA on the rest of the RULER benchmark at 4K context length. This covers much harder long context tasks, such as multi-key retrieval (MK1,MK2,MK3), multi-query (MQ), multi-value (MV), variable tracking (VT), common word extraction (CWE), frequent word extraction (FWE), Hotpot-QA (HQA), and Squad-QA (SQA). We compare LoLA against a stronger version of LoLCATs—with an equivalent, larger cache size ($\eta=896$)—and Mamba2-8B (Dao & Gu, 2024; Waleffe et al., 2024).
Table 2: Extended RULER Benchmark on 4K context lengths. Compared to NIAH tasks, these require much stronger forms of memory and state tracking. For cache parameters, LoLA uses $\eta=128,\lambda=768$ and LoLCATs uses $\eta=896$.
Model MK1 MK2 MK3 MQ MV VT CWE FWE HQA SQA Avg
Mamba2-8B 40.3 13.8 5.5 49.1 35.0 76.5 32.9 76.6 31.8 35.5 39.7
LoLCATs-8B+ 12.8 1.4 0.4 3.3 3.6 0.7 3.0 13.6 14.2 14.0 6.7
LoLA-8B 39.4 11.6 7.6 67.6 65.0 85.2 45.9 51.3 24.2 53.9 45.2
LoLA improves recall with minimal additional caching. Table 1 demonstrates an improvement from the base model’s 0.6% to 97.4% accuracy on S-NIAH-1. This is achieved with a $4.6\times$ smaller cache than Llama with $\eta=256,\lambda=256$. Furthermore, we show in Table 2 that sparse caching is essential for more difficult tasks, improving an extended form of LoLCATs from 6.7% average accuracy to 45.2%. For example, tasks such as variable tracking (VT) require understanding all of the context. Since no part of the sequence can be lost, naive metrics for sparse attention, such as (Zhang et al., 2023), cannot be used. Memory collisions must be mitigated. Sparse caching—specifically with our self-recall error—unlocks a new capability for hybrid linear attention architectures.
@@ 4.2 Commonsense Reasoning
We demonstrate the language modeling performance of LoLA on various zero-shot commonsense reasoning tasks using LM evaluation harness (Gao et al., 2024). To compare against previously available approaches, we use PIQA (PI) (Bisk et al., 2020), ARC-Easy (AE) & ARC-Challenge (AC) (Clark et al., 2018), HellaSwag (HS) (Zellers et al., 2019), WinoGrande (WG) (Sakaguchi et al., 2021), MMLU (MM) (Hendrycks et al., 2020), and Lambada OpenAI (Paperno et al., 2016)).
In Table 3, we compare LoLA against other 7-9B subquadratic models (Mamba (Gu & Dao, 2024), Mamba2 (Dao & Gu, 2024), RWKV-6 (Peng et al., 2024), Hawk & Griffin (De et al., 2024), Falcon Mamba (Zuo et al., 2024), RecurrentGemma (Botev et al., 2024), Mamba-in-the-Llama (Wang et al., 2024), Llamba (Bick et al., 2025), Hedgehog (Zhang et al., 2024), and LoLCATs (Zhang et al., 2025a)). Models that use any form of unbounded global attention (e.g interleaving SSM blocks and Transformer blocks) still retain quadratic compute complexity and growing memory costs. These are outside the scope of this work. We also report the number of training tokens used to create each model in both tables. Though LoLA is a training-free inference strategy that can be used for any sliding window + linear attention model, we report the cost of distilling the base subquadratic model (Zhang et al., 2025a). In Appendix A, we show results for 1-2B subquadratic models. Additionally, we provide a direct comparison of distilled models by measuring the average accuracy relative to their teacher models.
Table 3: Performance comparison of 7-9B parameter fixed-memory models across various common sense reasoning tasks: PIQA (PI), Arc-Easy (AE), ARC-Challenge (AC), Winogrande (WG), MMLU ($M$), and Lambada-openai (LB). Bolded scores are the best and underlined scores are the second best. We report accuracy for all applicable, except AC and HS use normalized logits. MMLU (MM) uses 5-shot. Reported scores were compiled from (Bick et al., 2025; Zhang et al., 2025a; Waleffe et al., 2024; Wang et al., 2024). * indicates our reproduced score is used and is higher than reported score. LoLA uses small cache parameters, $\eta=\lambda=64$.
Model Tokens (B) PI AE AC HS WG $M$ LB
@@
On short context tasks such as Winogrande, we observe additional caching is not needed as only local information is required for good performance. On the other hand, we find that gaps still exist with Lambada and MMLU. Lambada requires longer context reasoning, leading to significant improvements with sparse caching. Furthermore, Bick et al. (Bick et al., 2025) suggest that dataset selection plays a large role for MMLU performance. Though sparse caching shows significant improvement on MMLU, a more powerful distillation procedure or dataset may be needed to reach Llama’s performance (Goldstein et al., 2025).
@@ 4.3 Understanding Memory Collisions
We visualize how memory collisions occur in practice. At each time step $t$, we measure the self-recall error from equation 10 for every KV pair that is currently stored in the hidden state $\mathbf{H}_{t}$. We visualize the error for linear attention, sliding window + linear attention, and LoLA. We use a sliding window size of $\eta=64$ tokens and a sparse cache size of $\lambda=64$ when applicable.
Figure 3: Visualizing memory collisions by measuring SRE for stored KV pairs.
When only using linear attention, we observe large recall errors. At early time steps, the first few KV pairs receive small errors, but quickly become larger after hidden state updates. In Appendix G, we additionally plot the relative error to better show how this occurs. These stored associative memories become corrupted and tough to recall in the future. Furthermore, difficult-to-memorize pairs are evident, illustrated as bright columns in Figure 3. The additional use of sliding window attention only delays the inevitable memory collisions.
@@
In our final experiment, we measure alternative scoring functions to understand which KV pairs should be sparsely cached. Traditional sparse attention metrics assume "unimportant" tokens are evicted entirely from the context (Zhang et al., 2023; Singhania et al., 2024; Zaheer et al., 2020). In our setting, these assumptions are invalid as unimportant tokens are stored in low precision through linear attention.
Table 4: Ablation results for various scoring methods on S-NIAH-1 with 512 context length, $\eta=64,\lambda=64$. Extended details for the score calculation can be found in Appendix F.
Importance Metric S-NIAH-1 @ .5K Informal Assumption for “Important” Pairs $\left\|\cfrac{\phi(\bm{k})^{\top}\mathbf{H}}{\phi(\bm{k})^{\top}\bm{s}}-\bm{v}\right\|$ 99.0% Pairs that do not align with the hidden state’s prediction $(\exp(\bm{q}^{\top}\bm{k})-\phi(\bm{q})^{\top}\phi(\bm{k}))^{2}$ 11.4% Keys with incorrect attention weights $|\exp(\bm{q}^{\top}\bm{k})-\phi(\bm{q})^{\top}\phi(\bm{k})|$ 20.0% Keys with incorrect attention weights $\cfrac{\phi(\bm{q})^{\top}\phi(\bm{k})}{\exp(\bm{q}^{\top}\bm{k})}$ 52.0% Keys that Linear Attention over estimates $\exp(\bm{q}^{\top}\bm{k})$ 10.6% Keys that are attended to during sliding window attention None, extend sliding window 29.4% Most recent pairs
In Table 4, we observe that storing keys with poor exponential dot product approximations underperforms the naive extension of sliding window attention. This hints that a better softmax approximation should not be the main objective for linear attention methods. Keys over-estimated by linear attention seem to be “more important” than local keys; however, all tested alternatives fall short of enabling associative recall.
We also compare against traditional sparse attention ideas that use softmax attention scores as a proxy for importance, such as in H2O (Zhang et al., 2023) and LESS (Dong et al., 2024). Specifically, we found that a key’s average similarity score, $\exp(\bm{q}^{\top}\bm{k})$, does not translate well in the hybrid linear attention setting. Since queries are used to influence the selection of keys in the sparse cache, we believe highly similar phrases in the haystack may interfere with finding the needle. LoLA, on the other hand, benefits from its query-agnostic metric.
@@ 5 Conclusion
LoLA integrates linear attention with sparse caching to effectively mitigate memory collisions. By selectively retaining KV pairs that do not align with the current hidden state, LoLA enables passkey retrieval when the base model fails. Our experimental results demonstrate that targeted sparse caching substantially improves long context performance over naively increasing the sliding window size. LoLA demonstrates strong language modeling performance over other 1B or 8B subquadratic models.
Future work. The sparse cache carries a small overhead compute cost of $\mathcal{O}(\lambda d)$ for scoring. For high-complexity, long-context tasks, we found that larger sparse cache sizes are needed to reduce interference in the hidden state. We believe that these limitations can addressed in the future with a better base subquadratic model. More advanced base architectures, such as LaCT (Zhang et al., 2025b) or Atlas (Behrouz et al., 2025), use nonlinear key-to-value maps, which may lead to smaller caches.
Reproducibility Statement
For reproducibility of distilling base subquadratic model, see (Zhang et al., 2025a). This also includes released weights for the 8B parameter model on Huggingface. All evaluations were performed using LM Evaluation Harness (Gao et al., 2024). pseudo-code is available for reproducing LoLA’s attention operation in Appendix H. This is a drop-in replacement for the hybrid linear attention in the base model. Full code will be provided in the camera-ready version. Lastly, LLMs had minor contributions to the paper writing, such as spell-check and formatting.
@@ Appendix A Extended Language Modeling Results
Following Section 4.2, we extend this comparison in Table 5 for various 1-2B parameter subquadratic models (Grattafiori et al., 2024; Li et al., 2023; Sun et al., 2023; Qin et al., 2024; Schlag et al., 2021; Yang et al., 2024a; Gu & Dao, 2024; Dao & Gu, 2024; Beck et al., 2024; Peng et al., 2024; Bick et al., 2024; Zhang et al., 2025a; Bick et al., 2025; Ren et al., 2025). Compared to the 8B models, we observe that sparse caching is more important in the 1B parameter regime since there exist more memory collisions. This is a direct result of a smaller hidden state dimension. The size of $\mathbf{H}_{t}$ scales with the head dimension of the base model. Llama-3.2 1B’s head dimension is half that of Llama-3.1 8B.
@@
Subquadratic: Pretrained from scratch on FineWeb-Edu (Yang et al., 2024a; Penedo et al., 2024)
@@
Subquadratic: Pretrained from scratch on various sources (Bick et al., 2024; 2025)
@@
Subquadratic: Distilled from Phi-1.5-1.3B (Bick et al., 2024; Zhang et al., 2025a)
@@
Subquadratic: Distilled from Llama-3.2-1B (Bick et al., 2025; Zhang et al., 2025a)
@@
Llamba-1.3B 8 74.0* 69.5* 37.2 61.2 60.6 48.4 58.5 $0.975\times$
LoLCATs-1.3B 0.04 74.6 63.0 35.1 63.7 61.5 53.4 58.6 $0.977\times$
LoLA-1.3B (ours) 0.04 76.2 66.2 36.9 64.1 60.9 61.9 61.0 $1.017\times$
@@
Mamba2-Llama3-8B 20 76.8 74.1 48.0 70.8 58.6 43.2 61.9 $0.837\times$
@@
Llamba-8B 12 80.9 82.5 54.6 77.6 73.3 69.4 73.1 $0.987\times$
LoLCATs-8B 0.04 81.0 82.4 54.4 79.1 73.6 67.6 73.0 $0.987\times$
LoLA-8B 0.04 81.6 82.5 55.4 79.8 73.6 74.9 74.6 $1.009\times$
@@
Appendix B Efficiency Analysis
Cache Parameters.
Here, we analyze the efficiency of chunkwise LoLA. For various sliding window ($\eta$) and sparse cache ($\lambda$) sizes, we measure the peak VRAM cost and Time-to-First-Token for LoLA-8B with 4K long context on an Nvidia RTX 4090. Additionally, we show how different cache parameters lead to varying performance on RULER’s variable tracking task (Hsieh et al., 2024). There exists a trade-off between speed, memory footprint, and performance. There is no one-size-fits-all solution; we provide a short guide on navigating this tradeoff.
Figure 4: Measuring Time-to-First-Token for various sliding window and sparse cache sizes. This measurement is averaged across 100 trials and assumes data is already loaded into VRAM.
Figure 4 illustrates that optimal throughput is achieved by maximizing the sliding window size that fits into VRAM and minimizing the sparse cache size. This reduces the number of chunks computed in sequential order, allowing for more intra-chunk parallelization.
Figure 5: Measuring Peak VRAM usage for various sliding window and sparse cache sizes. This measurement includes the base model weights, the data sequence, and online activations such as KV caches.
Figure 5 suggests the total cache size ($\eta$ & $\lambda)$ needs to be reduced to lower VRAM cost. In our implementation, the data (4K long context sequence) already exists in VRAM, so this is included in the peak VRAM measurement. Furthermore, the TTFT does not include loading this data.
Figure 6: Cache Sizes vs. Variable Tracking performance at 4K context length from RULER (Hsieh et al., 2024)
Lastly, we show the variable tracking performance for various cache parameters in Figure 6. Variable tracking requires understanding all of the context. The base subquadratic model—or even extended sliding window variants—cannot perform this task without a sparse cache. As a guideline to increase general long context performance, the sparse cache size should be maximized. This mitigates memory collisions, preserving linear attention’s hidden state on long sequences.
In summary, LoLA introduces a new trade-off for subquadratic models. Low VRAM and high performance can be achieved (maximize $\lambda$, minimize $\eta$), but the model will be slow. Low VRAM, fast models (minimize $\lambda$, moderate $\eta$) will not be able to perform well on long-context tasks (e.g. the base subquadratic model, LoLCATs). Finally, fast and high performing models will require a much larger memory footprint (maximize both $\lambda$ and $\eta$). This will extend the applicability of subquadratic models across various hardware platforms such as small inference chips or large training servers.
LoLA Cache vs. Transformers.
In Figure 7, we compare LoLA’s bounded inference costs with vanilla, softmax attention. We compute the cache size as the total number of elements in all vectors and matrices stored for each attention head. For transformers, we store key and value vectors for each token in context, $t\,(d_{k}+d_{v})$. For LoLA, we add up the elements in each of the three memory systems: sliding window cache, sparse cache, and linear attention’s hidden state & normalizing state. Here, we provide both "small" and "large" cache size variants of LoLA for reference. In short context lengths, LoLA does not instantiate linear attention’s hidden state or measure the SRE. LoLA is equally efficient to softmax attention here.
Figure 7: Cache Size vs. Context Length for LoLA and vanilla transformers.
This figure illustrates why we are interested in subquadratic models to begin with. As context scales, we must linearly increase VRAM and compute per token. For example, we observed out-of-memory errors with Llama-3.1-8B at 4K context length in our experiments. LoLA, on the other hand, can be scaled with $\eta$ and $\lambda$ to maximize performance on specific hardware. The VRAM cost is agnostic to context length, meaning this model will always be able to fit.
Appendix C Extended Long Context Tasks
To further extend our needle-in-a-haystack results from Table 1, we provide more scores in Table 7 across a greater variety of cache parameter combinations. For simplicity, we chose $\eta=\lambda$ and varied the total cache size, marked with "++". Each additional "++" doubles the total cache size (i.e baseline holds 64 tokens, + holds 128, ++ holds 256, etc.). Additionally, we provide longer sequences for S-NIAH-1 in Table 8.
In these tasks, the “haystack” is synthetically constructed with various context lengths. The first task, S-NIAH-1, uses random sentences (e.g., “The grass is green.“) as the haystack, while S-NIAH-2 & 3 use essays. The needle—represented as a (word, number) pair—is placed in the haystack. At the end of the prompt, the model is tasked with returning the associated number with the special word. The first two tasks (S-NIAH-1 & 2) use a 7-digit number in the needle, and S-NIAH-3 uses a 32-digit UUID, requiring more tokens to represent the needle.
[…]
Model .5K 1K 2K 4K 8K 16K
[…]
In general, extending the cache size can improve performance while still maintaining high compression rates over transformers. For even longer context lengths, LoLA’s cache can easily be scaled at inference to achieve the desired recall performance. This can be seen in Table 8, where LoLA performs well up to 8K context length, which is $8\times$ longer than the sequences seen during distillation.
For large haystacks, the accuracy of LoLCATs is roughly the proportion of context that the sliding window covers. For smaller haystacks, the performance is slightly higher than that ratio since fewer pairs are stored in the hidden state. This results in fewer collisions; though of course, this does not scale. Additionally, we observe that fewer collisions exist in essay-based haystacks (S-NIAH-2 & 3), likely as a result of being more similar to the distillation data. Lastly, we observe that harder needle-in-a-haystack tasks (e.g., S-NIAH-3 with 32-digit needles) may require more sparse caching. To further extend these results, we believe training with longer sequences should yield stronger performance.
Appendix D Related Work